### v1

@@ Abstract
Diffusion models have emerged as a dominant approach for text-to-image generation. Key components such as the human preference alignment and classifier-free guidance play a crucial role in ensuring generation quality. However, their independent application in current text-to-image models continues to face significant challenges in achieving strong text-image alignment, high generation quality, and consistency with human aesthetic standards. In this work, we for the first time, explore facilitating the collaboration of human performance alignment and test-time sampling to unlock the potential of text-to-image models. Consequently, we introduce CHATS (Combining Human-Aligned optimization and Test-time Sampling), a novel generative framework that separately models the preferred and dispreferred distributions and employs a proxy-prompt-based sampling strategy to utilize the useful information contained in both distributions. We observe that CHATS exhibits exceptional data efficiency, achieving strong performance with only a small, high-quality funetuning dataset. Extensive experiments demonstrate that CHATS surpasses traditional preference alignment methods, setting new state-of-the-art across various standard benchmarks.
@@ 2.1 Human Preference Optimization
Human preference optimization aims to improve the aesthetic quality of generated images. One line of research Xu et al. (2024); Wu et al. (2023); Kirstain et al. (2023b); Lee et al. (2023) focused on developing reward models to evaluate the alignment between generated images and their corresponding prompts. These reward models are subsequently used to finetune diffusion models, guiding them to better capture desirable characteristics. Another line of research explored direct finetuning on preferred images Dong et al. (2023); Dai et al. (2024) or adopted reinforcement learning from human feedback (RLHF) Black et al. (2024); Fan et al. (2023); Yang et al. (2024); Wallace et al. (2024). Specifically, DDPO Black et al. (2024) formulated denoising as a multi-step decision-making problem. DPOK Fan et al. (2023) incorporated KL regularization as an implicit reward to stabilize the finetuning process. D3PO Yang et al. (2024) and Diffusion-DPO Wallace et al. (2024) extended the concept of DPO Rafailov et al. (2023) to diffusion models by optimizing the respective policies. Different from these methods, CHATS employs two distinct models to separately capture preferred and dispreferred information.
@@ 4.2 Sampling
In this formulation, $p_{\theta^{+}}(z_{t}|c)^{1+s}$ encourages the samples to align closely with the preferred distribution. When $0<\alpha<\frac{1+s}{s}$, the dispreferred distribution $p_{\theta^{-}}$ is partially incorporated, allowing it to contribute potentially useful patterns while remaining less influential than $p_{\theta^{+}}$. Conversely, when $\alpha<0$, the terms $p_{\theta^{-}}(z_{t}|c)^{\alpha s}$ and $p_{\theta^{-}}(z_{t})^{-(1+\alpha)s}$ actively push the samples away from undesired modes in the dispreferred distribution, both conditionally and unconditionally.

### v2

@@ Abstract
Diffusion models have emerged as a dominant approach for text-to-image generation. Key components such as the human preference alignment and classifier-free guidance play a crucial role in ensuring generation quality. However, their independent application in current text-to-image models continues to face significant challenges in achieving strong text-image alignment, high generation quality, and consistency with human aesthetic standards. In this work, we for the first time, explore facilitating the collaboration of human performance alignment and test-time sampling to unlock the potential of text-to-image models. Consequently, we introduce CHATS (Combining Human-Aligned optimization and Test-time Sampling), a novel generative framework that separately models the preferred and dispreferred distributions and employs a proxy-prompt-based sampling strategy to utilize the useful information contained in both distributions. We observe that CHATS exhibits exceptional data efficiency, achieving strong performance with only a small, high-quality funetuning dataset. Extensive experiments demonstrate that CHATS surpasses traditional preference alignment methods, setting new state-of-the-art across various standard benchmarks. The code is publicly available at github.com/AIDC-AI/CHATS.
@@ 2.1 Human Preference Optimization
Human preference optimization aims to improve the aesthetic quality of generated images. One line of research Xu et al. (2024); Wu et al. (2023); Kirstain et al. (2023b); Lee et al. (2023) focused on developing reward models to evaluate the alignment between generated images and their corresponding prompts. These reward models are subsequently used to finetune diffusion models, guiding them to better capture desirable characteristics.
Another line of research explored direct finetuning on preferred images Dong et al. (2023); Dai et al. (2024) or adopted reinforcement learning from human feedback (RLHF) Black et al. (2024); Fan et al. (2023); Yang et al. (2024); Wallace et al. (2024). Specifically, DDPO Black et al. (2024) formulated denoising as a multi-step decision-making problem. DPOK Fan et al. (2023) incorporated KL regularization as an implicit reward to stabilize the finetuning process. D3PO Yang et al. (2024) and Diffusion-DPO Wallace et al. (2024) extended the concept of DPO Rafailov et al. (2023) to diffusion models by optimizing the respective policies. Different from these methods, CHATS employs two distinct models to separately capture preferred and dispreferred information.
@@ 4.2 Sampling
Please refer to Sec A.4 for the details of its theoretical foundations. In this formulation, $p_{\theta^{+}}(z_{t}|c)^{1+s}$ encourages the samples to align closely with the preferred distribution. When $0<\alpha<\frac{1+s}{s}$, the dispreferred distribution $p_{\theta^{-}}$ is partially incorporated, allowing it to contribute potentially useful patterns while remaining less influential than $p_{\theta^{+}}$. Conversely, when $\alpha<0$, the terms $p_{\theta^{-}}(z_{t}|c)^{\alpha s}$ and $p_{\theta^{-}}(z_{t})^{-(1+\alpha)s}$ actively push the samples away from undesired modes in the dispreferred distribution, both conditionally and unconditionally.
@@ A.2 Training Objective of CHATS
A.3 Convergence Properties of CHATS
Given that DPO is invariant to affine transformations of the reward, for reward: $R^{\prime}(z_{0:T},c)=a\cdot R(z_{0:T},c)+b$, the optimal policy becomes (cf. Eq. 29):
$p^{*}(z_{0:T}\mid c)=\frac{p_{\mathrm{ref}}(z_{0:T}\mid c)e^{a\cdot R(z_{0:T},c)+b}}{Z^{\prime}(c)}.$ (45)
CHATS decomposes reward of traditional DPO (Eq. 30) into two parts:
$R^{+}(\theta^{+})=\log\frac{p_{\theta^{+}}(z_{0:T}^{+}\mid c)}{p_{\mathrm{ref}}(z_{0:T}^{+}\mid c)},\quad R^{-}(\theta^{-})=\log\frac{p_{\theta^{-}}(z_{0:T}^{-}\mid c)}{p_{\mathrm{ref}}(z_{0:T}^{-}\mid c)},$ (46)
with $\beta$ and $\log Z(c)$ omitted since they are constants for optimization. Defining the effective reward as: $R_{\mathrm{CHATS}}=R^{+}(\theta^{+})+R^{-}(\theta^{-})$, the resulting joint optimal distribution becomes:
$p^{*}(z_{0:T}\mid c)=\frac{p_{\mathrm{ref}}(z_{0:T}\mid c)e^{R_{\mathrm{CHATS}}}}{Z^{\text{CHATS}}(c)}.$ (47)
Under the assumption of $L$-smoothness and using standard gradient descent, we obtain the recursive inequality:
$\mathcal{L}_{k+1}\leq\mathcal{L}_{k}-\frac{\eta}{2}\|\nabla\mathcal{L}(\theta^{+}_{k},\theta^{-}_{k})\|^{2},$ (48)
which ensures the CHATS loss (Eq. 9) decreases monotonically and converges. Since their combination recovers the same optimal joint distribution as traditional DPO methods Wallace et al. (2024), CHATS preserves theoretical foundations of DPO.
A.4 Theoretical Foundations of CHATS Sampler
Start from Bayes’ rule for a classifier:
$p(c\mid z_{t})=\frac{p(z_{t}\mid c)p(c)}{p(z_{t})},$ (49)
since $p(c)$ can be regarded as a constant during optimization, CFG defines the guided distribution by raising $p(c\mid z_{t})$ with a guidance scale $s$:
$\tilde{p}(z_{t}\mid c)\propto p(z_{t}|c)\bigl{[}p(c\mid z_{t})\bigr{]}^{s}.$ (50)
Substitute the expression for $p(c\mid z_{t})$ and omit $p(c)$ yields:
$\tilde{p}(z_{t}\mid c)\propto p(z_{t}\mid c)^{1+s}p(z_{t})^{-s}.$ (51)
In CHATS, two models are used:
  • The preferred model $p_{\theta^{+}}(z_{t}\mid c)$,
  • The dispreferred model $p_{\theta^{-}}(z_{t}\mid c)$ (with its unconditional form $p_{\theta^{-}}(z_{t})$).
For each model, we can write a classifier-like term via Bayes’ rule. For the preferred model:
$p_{\theta^{+}}(c\mid z_{t})=\frac{p_{\theta^{+}}(z_{t}\mid c)p(c)}{p_{\theta^{+}}(z_{t})},$ (52)
and for the dispreferred model:
$p_{\theta^{-}}(c\mid z_{t})=\frac{p_{\theta^{-}}(z_{t}\mid c)p(c)}{p_{\theta^{-}}(z_{t})}.$ (53)
Assuming $p_{\theta^{+}}(z_{t})\approx p_{\theta^{-}}(z_{t})$, we combine the two signals by defining a composite log-odds score:
$\Delta(z_{t},c)=\log\frac{p_{\theta^{+}}(z_{t}\mid c)}{p_{\theta^{-}}(z_{t})}+\alpha\log\frac{p_{\theta^{-}}(z_{t}\mid c)}{p_{\theta^{-}}(z_{t})}.$ (54)
The first term tends to generate features favored by the preferred model while suppressing the background features typically produced by the dispreferred model in its unconditional output (similar to CFG), and the second term further accounts for the shift in the output of the dispreferred model when conditioned on $c$, with its impact regulated by a scalar $\alpha$. In this form, the useful information in $p_{\theta^{-}}(z_{t}\mid c)$ is effectively utilized as well.
Following CFG, we define the CHATS guided distribution as:
$\tilde{p}_{\theta}(z_{t}\mid c)\propto p_{\theta^{+}}(z_{t}\mid c)\exp\Bigl{(}s\cdot\Delta(z_{t},c)\Bigr{)}.$ (55)
Substitute $\Delta(z_{t},c)$:
$\tilde{p}_{\theta}(z_{t}\mid c)\propto p_{\theta^{+}}(z_{t}\mid c)\cdot\exp\left(s\left[\log\frac{p_{\theta^{+}}(z_{t}\mid c)}{p_{\theta^{-}}(z_{t})}+\alpha\cdot\log\frac{p_{\theta^{-}}(z_{t}\mid c)}{p_{\theta^{-}}(z_{t})}\right]\right).$ (56)
Using $\exp(s\log A)=A^{s}$, we have
$\tilde{p}_{\theta}(z_{t}\mid c)\propto p_{\theta^{+}}(z_{t}\mid c)\left(\frac{p_{\theta^{+}}(z_{t}\mid c)}{p_{\theta^{-}}(z_{t})}\right)^{s}\left(\frac{p_{\theta^{-}}(z_{t}\mid c)}{p_{\theta^{-}}(z_{t})}\right)^{\alpha s}.$ (57)
Grouping terms, we obtain
$\tilde{p}_{\theta}(z_{t}\mid c)\propto p_{\theta^{+}}(z_{t}\mid c)^{1+s}p_{\theta^{-}}(z_{t}\mid c)^{\alpha s}p_{\theta^{-}}(z_{t})^{-(1+\alpha)s},$ (58)
which is the same with Eq. 15. The final guided distribution is not merely a sharpened version of $p_{\theta^{+}}(z_{t}\mid c)$. It also leverages the dispreferred model. The term$\left(\frac{p_{\theta^{-}}(z_{t}\mid c)}{p_{\theta^{-}}(z_{t})}\right)^{\alpha s}$ adjusts the output based on how conditioning on $c$ changes the dispreferred model’s behavior. This derivation, starting from $p(c\mid z_{t})$ for both models, provides a theoretical foundation for the CHATS sampling distribution analogous to that of CFG.